find class boundaries,midpoints,and widths for each class?
a.11-15
b.17-39
c.293-353
d.11.8-14.7
e.3.13-3.93
1
Expert's answer
2020-01-20T09:31:32-0500

a.11-15

Class boundaries = 11 - 0.5 and 15 + 0.5

So class boundaries = 10.5 and 15.5

Midpoint = (10.5 + 15.5)/2 = 13

Class width = 15.5 - 10.5 = 5


b.17-39

Class boundaries = 17 - 0.5 and 39 + 0.5

So class boundaries = 16.5 and 39.5

Midpoint = (16.5 + 39.5)/2 = 28

Class width = 39.5 - 16.5 = 23


c.293-353

Class boundaries = 293 - 0.5 and 353 + 0.5

So class boundaries = 292.5 and 353.5

Midpoint = (292.5 + 353.5)/2 = 323

Class width = 353.5 - 292.5 = 61


d.11.8-14.7

Class boundaries = 11.8 - 0.05 and 14.7 + 0.05

So class boundaries = 11.75 and 14.75

Midpoint = (11.75 + 14.75)/2 = 13.25

Class width = 14.75 - 11.75 = 3


e.3.13-3.93

Class boundaries = 3.13-0.005 and 3.93 + 0.005

So class boundaries = 3.125 and 3.935

Midpoint = (3.125 + 3.935)/2 = 3.53

Class width = 3.935 - 3.125 = 0.81